### v2

# M3D: Manifold-based Domain Adaptation with Dynamic Distribution for Non-Deep Transfer Learning in Cross-subject and Cross-session EEG-based Emotion Recognition

Abstract
Emotion decoding using Electroencephalography (EEG)-based affective brain-computer interfaces (aBCIs) is crucial for affective computing but is hindered by EEG’s non-stationarity, individual variability, and the high cost of large-scale labeled data. Deep learning-based approaches, while effective, require substantial computational resources and large datasets, limiting their practicality. To address these challenges, we propose Manifold-based Domain Adaptation with Dynamic Distribution (M3D), a lightweight non-deep transfer learning framework. M3D includes four main modules: manifold feature transformation, dynamic distribution alignment, classifier learning, and ensemble learning. The data undergoes a transformation onto an optimal Grassmann manifold space, enabling dynamic alignment of the source and target domains. This process prioritizes both marginal and conditional distributions according to their significance, ensuring enhanced adaptation efficiency across various types of data. In the classifier learning, the principle of structural risk minimization is integrated to develop robust classification models. This is complemented by dynamic distribution alignment, which refines the classifier iteratively. Additionally, the ensemble learning module aggregates the classifiers obtained at different stages of the optimization process, which leverages the diversity of the classifiers to enhance the overall prediction accuracy. The proposed M3D framework is evaluated on two benchmark EEG emotion recognition datasets using two validation protocols (cross-subject single-session and cross-subject cross-session), as well as on a clinical EEG dataset of Major Depressive Disorder (MDD). Experimental results demonstrate that M3D outperforms traditional non-deep learning methods, achieving an average improvement of 4.47%, while achieving deep learning-comparable performance with significantly lower data and computational requirements. These findings highlight the potential of M3D to enhance the practicality and applicability of aBCIs in real-world scenarios.
Index Terms:
Electroencephalography, Emotion Recognition, Dynamic Domain Adaptation, Non-Deep Transfer Learning, Manifold Transformation
11footnotetext:  Equal contributions.
I Introduction
Emotion, functioning as both a physiological and psychological state within an individual, serves as a rich repository of insights into mental and physical well-being [1, 2, 3]. Emotion recognition stands as a crucial method for objectively discerning human emotional states. Currently, the practice relies on the analysis of either non-physiological signals or physiological signals. In contrast to cues like facial expressions, voice intonation, and gestures, Electroencephalography (EEG) stands out as a non-invasive brain information collection technique [4]. Its strong correlation with human emotional states, coupled with its resistance to deception, positions EEG as a highly accurate means to objectively reflect human emotions [5, 6, 7].
EEG signals exhibit non-stationary characteristics and individual differences [8, 9], leading to significant variations in the distribution of EEG signals within the same subject across different time periods or between different individuals. Traditional machine learning methods assume independence and identical distribution of data, causing emotion recognition models to have poor generalization across different sessions and different individuals. This limitation poses a challenge in expanding the widespread adoption and application of these models to new sessions and individuals. On the other hand, EEG signals encounter the challenge of a small sample size, typically comprising only a few hundred to a few thousand samples [10]. The labeling process is often expensive and time-consuming, making it challenging to amass a large volume of data. This difficulty in obtaining substantial data poses a hurdle in training a reliable model.
Early EEG-based emotion recognition methods mainly rely on conventional non-deep machine learning algorithms[11], including Support Vector Machine (SVM), K-Nearest Neighbor (KNN), and Linear Discriminant Analysis (LDA). In these studies, there is an underlying assumption that EEG samples are independent and identically distributed. However, variations among individuals significantly undermine the effectiveness of traditional machine learning methods in cross-subject emotion recognition tasks. This results in suboptimal performance and poor generalization effects. On the other hand, transfer learning techniques can be beneficial in transferring informative data from the source (training data) to the target domain (test data), which could reduce the variations between different distributions in the modeling. More related transfer learning methods are detailed in Section II.
Compared to deep learning-based transfer learning methods, non-deep learning-based transfer learning offers a more lightweight approach. Unlike deep learning methods, non-deep learning-based transfer requires less labeled data, which is better suited to the current conditions in EEG tasks. In this paper, we propose a novel non-deep learning-based transfer learning framework for cross-subject emotion recognition, named as Manifold-based Domain adaptation with Dynamic Distribution (M3D). The M3D model maps EEG data onto a manifold space, where both marginal and conditional distributions are dynamically analyzed and aligned. This process is coupled with classifier learning to minimize distributional differences between source and target domains. Subsequently, ensemble learning integrates the results to enhance the robustness of cross-subject and cross-session emotion recognition. Experimental results demonstrate that the proposed M3D model achieves performance comparable to existing deep learning-based methods. The main contributions of the paper are summarized below.
  • A novel non-deep learning-based transfer learning model ((M3D) is proposed to address the variations in the distribution of EEG signals across different sessions and subjects.
  • Both marginal distribution and conditional distribution are considered during domain alignment, where a dynamically adjusted weighting of the importance of the two distributions is incorporated.
  • An adaptive classifier learning together with an ensemble learning is introduced to iteratively refine the classifier learning process and leverage the strength of multiple learning models to optimize the prediction of data labels.
  • Extensive experiments are conducted on two well-known databases using two different validation protocols (cross-subject single-session and cross-subject cross-session), as well as a clinical EEG database of Major Depressive Disorder (MDD). The results obtained from the proposed M3D model (a non-deep learning approach) are comparable to those achieved by existing deep learning methods.
Figure 1: The overall architecture of the proposed M3D model, which includes four mains modules. (1) Manifold Feature Transformation: incorporates TCA (Transfer Component Analysis) to obtain the embedded subspaces, and develops GFK (Geodesic Flow Kernel) algorithm to transform the original features into manifold features. (2) Dynamic Distribution Alignment: dynamically adjusts the importance of the marginal and conditional distributions by learning an adaptive factor $\mu$ during the alignment process. (3) Classifier Learning: summarizes structural risk minimization and distribution alignment to iteratively optimize the classifier. (4) Ensemble Learning: harmonizes the classification results from the iterative classifier learning process and generates a more reliable and robust classification performance. Here, $Source$ and $Target$ refer to the source and target data, respectively. $\mathbb{G}$ is the Grassmann manifold, projected by the geodesic function $\Psi(t)$ ($t\in[0,1]$). The points $\Psi(0)$ and $\Psi(1)$ correspond to the embedded representations of the source and target domains in $\mathbb{G}$. $x_{i,j}$ and $x_{k,l}$ are the samples from the source and target data,respectively. $z^{\infty}_{i,j,k,l}$ are the corresponding data representation in the manifold space by the geodesic transformation. $H^{\infty}_{K}$ denotes the reproducing kernel Hilbert space (RKHS), and $G$ is the computed geodesics kernel.
II RELATED WORK
Domain adaptation stands out as the predominant method in transfer learning, frequently employed to address differences between the distributions of source and target data. It targets the reduction of distribution disparities between the two domains and align the feature distributions, achieving the assumption that the source and target data is independently and identically distributed [12]. Existing domain adaptation-based transfer learning algorithms can be generally categorized into two broad categories: deep transfer learning and non-deep transfer learning.
II-A Deep Transfer Learning Methods
The current trend in the development of transfer learning methods has mainly focused on deep learning-based approaches. For example, Li et al. [13] pioneered the use of the Domain Adversarial Neural Network (DANN) in the study of EEG-based emotion recognition. The optimization objective was to establish a shared common feature representation that mitigates distribution differences between the source and target domains. Building upon the aligned feature representation achieved with DANN, Zhou et al. [14] introduced a prototypical representation-based pairwise learning framework, aiming to further enhance cross-subject emotion recognition performance. However, the current DANN network is limited by its exclusive focus on addressing marginal distribution differences in EEG data among different individuals, neglecting joint distribution differences. Considering that individual differences in EEG arise from joint distribution differences between different EEG signals, Li et al. [15] extended the DANN network by incorporating the Joint Domain Adaptation Network. This extension provides a more comprehensive approach that considers both marginal and joint distribution differences in deep transfer learning studies.
Despite the notable achievements of the deep transfer learning methods in emotion recognition tasks, there are certain limitations. Firstly, deep learning methods demand a substantial number of training samples. However, in EEG tasks, the available samples are typically limited to a few hundred or a few thousand. Insufficient samples would constrain performance, leading to overfitting and a decrease in generalization ability. Secondly, deep transfer learning algorithms heavily rely on a substantial amount of accurately labeled source data. If there is noticeable noise in the labels from the source domain, the recognition performance would significantly deteriorate [16]. However, acquiring a sufficient number of samples with precise label information is not only costly but also time-consuming. Thirdly, deep transfer learning methods exhibit high computational complexity, substantial computational requirements, and involve slow and complicated training processes. The reliance on high-end hardware facilities is a significant demand, posing challenges in practical application environments where such resources may not be readily available. In contrast, non-deep transfer learning models exhibit lower complexity, reduced hardware requirements, and lightweight properties, showing the potential for good performance in widespread applications in real-life scenarios.
II-B Non-Deep Transfer Learning Methods
Non-deep transfer learning methods align the feature distribution between source and target domains using conventional machine learning approaches. For example, Pan et al. [17] introduced a Transfer Component Analysis (TCA) algorithm, leveraging the reproducing kernel Hilbert space to learn transfer information. Through maximizing the Mean Discrepancy (MMD), it successfully mitigated the marginal distribution differences. Zheng et al. [18] introduced a Transductive Parameter Transfer (TPT) algorithm to build personalized classifiers for subjects from various source domains. This algorithm mapped the individualized classifier parameters to the target domain, minimizing the MMD between the source and target domains and adapting to marginal distribution. Wang et al. [19] introduced a Stratified Transfer Learning (STL) algorithm, which constructed a weak classifier with source data and making predictions on the target domain. Based on this foundation, it further conducted a more thorough exploration of intra-class relationships, leveraging class-relatedness to accomplish adaptive spatial dimensionality reduction.
Although the transfer learning strategies mentioned above effectively mitigate differences between the distributions of source and target data by considering marginal or conditional distributions, it’s crucial to note that the actual disparities in EEG between subjects primarily exist in their joint distribution [15]. Current domain adaptation algorithms either focus on marginal or conditional distributions, neglecting to assess the combined contributions in domain adaptation process. It would lead to inefficient data utilization and suboptimal results. To tackle this issue, He [15] proposed a Joint Distribution Adaptation (JDA), which involved matching these two distributions with equal weights to approximate the adjustment of the joint distribution [20]. Building upon the proposed JDA, other studies extended its capabilities by incorporating regularization [21], sparse representation [22], structural consistency [23], and domain-invariant clustering [24].
However, many existing methods assume equal contributions from the conditional and marginal probability distributions, which is often unrealistic in real-world applications. The relative importance of these distributions in shaping the joint probability distribution varies based on domain similarity. For example, when the source and target domains are highly dissimilar, aligning the marginal distribution is more critical. Conversely, when the two domains exhibit high similarity, prioritizing conditional distribution alignment is more beneficial. Therefore, for better alignment between the two domains, dynamically adjusting the weighting of marginal and conditional distributions based on the actual data characteristics is a critical issue in non-deep transfer learning research.
Moreover, most existing approaches attempt to align source and target distributions directly within the original feature space. However, this approach can be ineffective, as certain features may be distorted due to domain shifts or may be inherently domain-specific. To address this issue, several approaches leverage subspace geometrical structures, particularly the Grassmann manifold, for subspace learning. These methods aim to transform data into a new projection space where domain-invariant features can be extracted, minimizing distributional differences between source and target domains [25, 26, 27]. This highlights the necessity of feature transformation prior to distribution alignment.
To address these critical challenges and enhance non-deep transfer learning performance in EEG-based emotion recognition, we propose the Manifold-based Domain adaptation with Dynamic Distribution (M3D) model. Based on the extracted differential entropy (DE) features from raw EEG signals, we first incorporate the TCA algorithm to reduce feature dimensionality and narrow the distribution gap between the features derived from the source and target domains. Subsequently, the features are embedded into the Grassmann manifold space through the computation of the Geodesic Flow Kernel (GFK)[28]. To enhance the adaptability of our framework, we introduce an adaptive factor that dynamically assesses the significance of both marginal and conditional distributions in the two domains. This factor aids in optimizing the embedding process and ensuring the effective representation of manifold features. Upon achieving a robust manifold feature representation, we proceed with classifier learning to optimize the classification performance under the complex conditions of cross-subject and cross-session scenarios. In this phase, conventional machine learning methods such as SVM, KNN, AdaBoost, Gaussian Naive Bayesian (GNB), Decision Tree (DT), and Bagging classifiers are employed as the initial classifiers. These classifiers are chosen due to their simplicity, computational efficiency, and strong adaptability to small datasets. Finally, the classification results derived from diverse parameter settings are harmonized through an ensemble learning approach. This ensemble strategy enhances the overall robustness and reliability of the proposed framework, ultimately contributing to more accurate and reliable classification outcomes.
III METHODOLOGY
In this section, we will provide a detailed introduction to the proposed M3D model. As shown in Fig. 1, M3D includes four main modules: manifold feature transformation, dynamic distribution alignment, classifier learning, and ensemble learning. In the manifold feature transformation module, the input data initially undergoes a reduction to lower-dimensional features through TCA. Then, a manifold kernel $G$ is learned to facilitate the transformation of the reduced features into an optimal manifold space. In the dynamic distribution alignment module, Structural Risk Minimization (SRM) [29] is incorporated to dynamically align the feature distribution within the obtained manifold space. This adaptive alignment process enhances the model’s capacity to effectively capture and adapt to variations in the underlying data distribution. In the classifier learning module, a classifier is constructed using the aligned manifold features. Initially, conventional machine learning methods are employed, and the classifier is then further iteratively optimized by solving the given classification loss. In the ensemble learning module, all the optimized classification results from the iterative process in the classifier learning module are harmonized to generate a more robust and reliable performance. These are the final results for cross-subject and cross-session EEG-based emotion recognition.
III-A Manifold Feature Transformation
Traditional feature alignment methods typically rely on original data features, which can be problematic due to deformations within the original feature space. These deformations make it difficult to effectively reduce differences between the source and target domains, negatively impacting model performance. Alternatively, the manifold space offers a solution by capturing the core aspects of data and representing original information in a more compact form [30, 31]. Importantly, features embedded in the manifold space often exhibit favorable geometric properties, enhancing their effectiveness in domain adaptation [25, 26, 27]. In the present study, our emphasis centers on leveraging the Grassmann manifold space to encapsulate the intrinsic essence of the data, which enables the creation of a more robust and dimensionality-reduced representation and offers a more effective way to align features across different domains.
In this paper, we introduce a manifold feature transformation process, as illustrated in Fig. 1. The source and target domains are represented as $D_{S}=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$ and $D_{T}=\{(x_{n+1},y_{n+1}),...,(x_{n+m},y_{n+m})\}$, respectively, with each domain featuring $D$-dimensional data representations. Through TCA[17], we map both domains into a lower $d$-dimensional feature space ($d\ll D$), denoted as $T_{S}$ and $T_{T}$. This process effectively reduces the dimensionality of the feature space, facilitating more efficient data analysis and interpretation. Subsequently, we embed the mapped $d$-dimensional feature subspace into the Grassmann manifold, denoted as $\mathbb{G}$. In this manifold space, our goal is to enhance the similarity between the source and target domains’ data as much as possible. Simultaneously, we aim to retain the unique attributes and inherent characteristics of each domain’s data. This approach ensures that, despite the alignment of the domains in a shared feature space, the distinctiveness and originality of each domain are preserved, maintaining the integrity and context of the data from both domains.
To make the data distributions of the two domains more similar in the Grassmann manifold, we either bring the source domain closer to the target domain or vice versa. Here, we construct geodesics within $\mathbb{G}$, treating the source and target domains as points in this manifold space. By navigating along these geodesics, we facilitate a systematic journey from the source domain towards the target domain, enabling a gradual harmonization of their data distributions within the Grassmann manifold and ensuring a smoother and more effective domain alignment. In the process of constructing geodesics, we apply the standard Euclidean metric to the Riemannian manifold. This allows the geodesic to be described by a parameterization function, $\Psi$, where $t$ ranges from 0 to 1, mapping to $\Psi(t)$ within the Grassmann manifold. Specifically, $\Psi(0)$ and $\Psi(1)$ correspond to the embedded representations of the source and target domains in $\mathbb{G}$, respectively. Thus, the geodesic delineates the trajectory by which $\Psi(0)$ evolves into $\Psi(1)$, representing the most direct path within the manifold space that connects these two domain representations. To ensure a smooth and continuous transformation across the manifold, the intermediate values of $t$ ($t\in(0,1)$), $\Psi(t)$ satisfies the following conditions:
$\Psi\left(t\right)=T_{S}U_{1}\Gamma\left(t\right)-R_{S}U_{2}\Sigma\left(t\right),$ (1)
where $T_{S}\in\mathbb{R}^{M\times d}$ denote the set of orthonormal base for the source domain. $M$ is the sample size. $R_{S}\in\mathbb{R}^{M\times(D-d)}$ signifies the orthogonal complement of $T_{S}$, serving to encapsulate the dimensions outside the primary subspace spanned by $T_{S}$. $U_{1}\in\mathbb{R}^{d\times d}$ and $U_{2}\in\mathbb{R}^{(D-d)\times d}$ are both orthogonal matrices, which can be obtained through the following two sets of Singular Value Decomposition (SVD) as
$T^{\top}_{S}T_{T}=U_{1}\Gamma V^{\top},R^{\top}_{S}T_{T}=-U_{2}\Sigma V^{\top}.$ (2)
Here, $T_{T}\in\mathbb{R}^{M\times d}$ refer to the set of orthonormal base for the target domain. $\Gamma$ and $\Sigma$ are both $d\times d$ diagonal matrices, where their diagonal elements being $\cos{\theta_{k}}$ and $\sin{\theta_{k}}$ ($k=1,2,...,d$), respectively. The variable $\theta_{k}$ denotes the angles between the orthonormal bases $T_{S}$ and $T_{T}$, with the constraint $0\leq\theta_{1}\leq\theta_{2}\leq...\leq\theta_{d}\leq\dfrac{\pi}{2}$. $\theta_{k}$ serve as a metric for quantifying the alignment between the domains within the $\mathbb{G}$, indicating how closely the subspaces represented by $T_{S}$ and $T_{T}$ are oriented to one another.
The operation $\Psi(t)^{\top}x$ represents the projection of a feature vector $x$ from the $d$-dimensional feature space onto the Grassmann manifold space $\mathbb{G}$, where $x$ could be any data sample originating from either the source or the target domain. The selection of an appropriate value or range of values for $t$ becomes a significant challenge, as it directly influences the quality of the projection for $x$. The optimal choice of $t$ ensures that the projected data maintains a balance between adapting to the new domain and preserving its original domain characteristics. This balance is essential for effective domain adaptation, as it allows for the integration of domain-specific features while minimizing the loss of critical information during the transition.
Here, we introduce an integration strategy to identify the optimal $t$ or a set of $t$ points between any two given data samples $x_{i}$ and $x_{j}$, and utilize a geodesic kernel to facilitate the mapping process. This methodology effectively transforms the $d$-dimensional feature space into an infinite-dimensional feature space, which serves to mitigate domain drift (a common challenge in domain adaptation). For each pair of data samples, $x_{i}$ and $x_{j}$, we iteratively calculate their projections onto $\Psi(t)$, where $t$ varies incrementally from 0 to 1. Through this process, each data sample is mapped across a continuum within the Grassmann manifold space, capturing the evolution of its feature representation as it transitions from being aligned with the source domain ($t=0$) to being aligned with the target domain ($t=1$). The culmination of this iterative projection process results in the creation of two infinite-dimensional feature vectors, denoted as $z^{\infty}_{i}$ and $z^{\infty}_{j}$,
[…]
where $z^{\infty}_{i}$ and $z^{\infty}_{j}$ encapsulate the comprehensive trajectory of each data sample across the manifold, providing a rich, multidimensional perspective on the domain adaptation process. This strategy enhances the adaptability and accuracy of domain adaptation by leveraging the inherent structure of the data within the manifold, thereby offering a more effective approach to overcoming the challenges of domain drift. From [28], the geodesic kernel is then determined as the inner product of the two infinite-dimensional vectors, as
[…]
where $G\in\mathbb{R}^{D\times D}$ is a positive semi-definite matrix, computed as
$G=\begin{bmatrix}T_{S}U_{1}&R_{S}U_{2}\end{bmatrix}\begin{bmatrix}\Lambda_{1}&\Lambda_{2}\\ \Lambda_{2}&\Lambda_{3}\end{bmatrix}\begin{bmatrix}U^{\top}_{1}&T^{\top}_{S}\\ U^{\top}_{2}&R^{\top}_{S}\end{bmatrix}.$ (6)
Here, $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ are all diagonal matrices, and the corresponding diagonal elements are given as
$\begin{split}\begin{cases}\lambda_{1,k}=1+\dfrac{\sin{2\theta_{k}}}{2\theta_{k}},\\ \lambda_{2,k}=\dfrac{\cos{2\theta_{k}}-1}{2\theta_{k}},\\ \lambda_{3,k}=1-\dfrac{\sin{2\theta_{k}}}{2\theta_{k}},\\ \end{cases}\end{split}$ (7)
where $k$ is in the range of $[1,d]$. The computed geodesics kernel $G$ enables the effective transformation of a data sample $x$, initially positioned within the $d$-dimensional feature space, into the Grassmann manifold space, as
$z=\sqrt{G}x.$ (8)
This transformed representation, $z$, integrates the intrinsic features of $x$ as it is projected through the manifold, leveraging the geometrical properties of $\mathbb{G}$ to achieve a more domain-adaptive feature representation and address domain drift in transfer learning. In the implement, we use DE features for manifold feature transformation across all databases.
III-B Dynamic Distribution Alignment
After the manifold feature transformation, the feature distributions between the source and target domains would exhibit greater similarity than seen with the original feature representation. Still, differences remain in both the marginal probability distribution and conditional probability distribution. Current distribution adaptation methods commonly assume equal importance for both the marginal distribution $P$ and conditional distribution $Q$ [32, 33]. Traditional distribution adaptation methods often treat the adaptation of marginal and conditional distributions as equally important, an assumption that may not always be applicable. In scenarios where significant differences exist between the source and target domains, the adaptation of the marginal distribution takes on increased importance. Conversely, in situations where the source and target domains are more closely aligned, the adaptation of the conditional distribution becomes more critical. Therefore, recognizing and adjusting the emphasis on marginal or conditional distribution adaptation based on the domains’ specific characteristics is essential for achieving successful distribution adaptation.
In the present study, an adaptive factor $\mu\in[0,1]$ is incorporated to dynamically adjust the emphasis placed on marginal and conditional distributions. This adaptive factor allows for the flexible allocation of importance between the two types of distributions based on the characteristics of the source and target domains, which is defined as
$\begin{split}\overline{D}_{f}\left(T_{S},T_{T}\right)=\left(1-\mu\right)D_{f}\left(P_{s},P_{t}\right)\\ +\mu\sum^{C}_{c=1}D_{f}^{(c)}\left(Q_{s},Q_{t}\right),\end{split}$ (9)
where $\overline{D}_{f}\left(T_{S},T_{T}\right)$ represents the dynamic distribution adaptation between the source and target domains within the manifold feature space. $D_{f}(P_{s},P_{t})$ is the marginal distribution adaptation, and $D_{f}^{(c)}(Q_{s},Q_{t})$ is the conditional distribution adaptation for the class $c$ ($c\in\{1,...,C\}$). The total number of classes is represented by $C$. When $\mu$ is close to 0, it indicates significant differences between the source and target domains, thus prioritizing the adaptation of marginal distributions. When $\mu$ approaches 1, it suggests a greater similarity between the domains, thereby emphasizing the need for conditional distribution adaptation. At the midpoint, $\mu=0.5$, the model treats both marginal and conditional distributions with equal importance.
In (9), the marginal distribution adaptation $D_{f}(P_{s},P_{t})$ is defined as
$D_{f}\left(P_{s},P_{t}\right)=\|E\left[f\left(z_{s}\right)\right]-E\left[f\left(z_{t}\right)\right]\|^{2}_{H_{K}},$ (10)
where $z_{s}$ and $z_{t}$ are the source and target data represented in the manifold feature space, given in (8). $f$ refers to the classifier. The term $H_{K}$ denotes the reproducing kernel Hilbert space (RKHS), which is a space of functions generated by the feature mapping $\Psi(\cdot)$. This mapping is critical for capturing the complex structures within the data by projecting it into a higher-dimensional space where linear separability is more feasible. For the computation of the conditional distribution adaptation $D_{f}^{(c)}(Q_{s},Q_{t})$, it calculate in a similar way, as
$D_{f}^{(c)}\left(Q_{s},Q_{t}\right)=\|E\left[f\left(z_{s}^{(c)}\right)\right]-E\left[f\left(z_{t}^{(c)}\right)\right]\|^{2}_{H_{K}},$ (11)
where $z_{s}^{(c)}$ and $z_{t}^{(c)}$ are the $c$-class source and target data represented in the manifold feature space (8). Then, (9) could be rewritten as
$\begin{split}\overline{D}_{f}\left(T_{S},T_{T}\right)=\left(1-\mu\right)\|E\left[f\left(z_{s}\right)\right]-E\left[f\left(z_{t}\right)\right]\|^{2}_{H_{K}}\\ +\mu\sum^{C}_{c=1}\|E\left[f\left(z_{s}^{(c)}\right)\right]-E\left[f\left(z_{t}^{(c)}\right)\right]\|^{2}_{H_{K}}.\end{split}$ (12)
Given the absence of label information for the target domain during the training phase, we cannot directly obtain $f(z_{t}^{(c)})$ in (12). To tackle this issue, we approximate $f(z_{t}^{(c)})$ using the class-conditional probability distribution [34]. Specifically, the process begins by training an initial weak classifier, such as a SVM or KNN, on the source data. This initial classifier, $f$, is subsequently utilized to infer pseudo-labels for the target data, serving as a provisional substitute for the unavailable true labels of the target domain. To enhance the reliability and accuracy of the pseudo-labels, an iterative refinement strategy is employed to improve the decodability $f$. This involves using the outcomes of the initial classifier to iteratively train subsequent classifiers, thereby progressively refining the pseudo-labels. This iterative process aims to converge towards more accurate pseudo-labels, thus improving the model’s ability to generalize from the source to the target domain effectively. Further details on the specific iterative refinement process and how it contributes to classifier learning will be discussed in Section III-C.
In (12), the adaptive factor $\mu$ is not a constant value predetermined by prior knowledge. Instead, it is dynamically learned based on the underlying data distribution. Specifically, we use the A-distance measurement [35] as a metric to estimate the marginal distribution difference, denoted as $d_{A}$, between the source and target domains, as
$d_{A}\left(T_{S},T_{T}\right)=2\left(1-2\epsilon\left(h\right)\right),$ (13)
where $\epsilon\left(h\right)$ denotes the hinge loss, which is derived from a binary classifier trained specifically to distinguish between samples from the source and target domains. For the conditional distribution difference, based on the provided label information for source data and the estimated pseudo-label information of target data, we conduct the A-distance measurement within each $c$ class ($c\in[1,\dots,C]$) as
$d_{c}=d_{A}\left(T_{S}^{(c)},T_{T}^{(c)}\right).$ (14)
Here, $T_{S}^{(c)}$ and $T_{T}^{(c)}$ represent the subsets of data corresponding to class $c$ from the source and target domains, respectively. Then, the adaptive factor $\mu$ can be estimated as
$\mu=1-\dfrac{d_{A}}{d_{A}+\sum^{C}_{c=1}d_{c}}.$ (15)
In the process of dynamic distribution alignment, the adaptive factor $\mu$ is recalculated at each iteration, underscoring the iterative and responsive nature of this adaptation strategy. This re-calibration is essential to ensuring that the adaptation process remains attuned to the evolving similarities and differences between the source and target domain feature distributions as they are represented in the manifold feature space.
III-C Classifier Learning
Through the integration of manifold feature learning and dynamic distribution alignment, we can formulate an adaptive classifier $f$ at the $\iota$th iteration as
$\begin{split}f^{\iota}=\mathop{\arg\min}_{f\in H_{K}}\sum^{n}_{i=1}\left(y_{i}-f^{(\iota-1)}\left(z_{i}\right)\right)^{2}+\eta\|f^{(\iota-1)}\|^{2}_{H_{K}}\\ +\lambda\overline{D}_{f^{(\iota-1)}}\left(T_{S},T_{T}\right)+\rho R_{f^{(\iota-1)}}\left(T_{S},T_{T}\right).\end{split}$ (16)
Here, $\iota$ is the iteration number, $y_{i}$ is the groundtruth, and $f(z_{i})$ is the predicted classification result. $\|\cdot\|^{2}_{H_{K}}$ refers to the L2 norm in the reproducing kernel Hilbert space. The dynamic distribution adaptation term, $\overline{D}_{f}\left(T_{S},T_{T}\right)$, is obtained in (12). Further, a Laplacian regularization term, $R_{f}\left(T_{S},T_{T}\right)$, is incorporated to leverage the inherent geometric similarities among neighboring points in the manifold $\mathbb{G}$ [30]. Here, the classifier $f$ can be initialized using a commonly employed machine learning classifier, followed by adaptive optimization through an iterative learning process.
In the SRM framework with the representer theorem [30], $f$ can be expanded as
$f\left(z\right)=\sum^{n+m}_{i=1}\beta_{i}K\left(z_{i},z\right),$ (17)
which includes both labeled and unlabeled samples from source and target domains, and $\beta=(\beta_{1},\beta_{2},\dots,\beta_{n+m})^{T}\in\mathbb{R}^{(n+m)\times 1}$ is the coefficient vector. $K$ is the kernel function. For the SRM on the source domain, the first part in (16) could be expressed as
$\begin{split}\sum^{n}_{i=1}\left(y_{i}-f\left(z_{i}\right)\right)^{2}+\eta\|f\|^{2}_{H_{K}}\\ =\sum^{n+m}_{i=1}A_{ii}\left(y_{i}-f\left(z_{i}\right)\right)^{2}+\eta\|f\|^{2}_{H_{K}},\\ \end{split}$ (18)
where $\|\cdot\|_{F}$ is the Frobenius norm. The matrix $A\in\mathbb{R}^{(n+m)\times(n+m)}$ is a diagonal indicator matrix with $A_{ii}=1$ if $i\in T_{S}$, and $A_{ii}=0$ if $i\in T_{T}$. Substituting (17) into (18), then we can obtain
$\begin{split}\sum^{n}_{i=1}\left(y_{i}-f\left(z_{i}\right)\right)^{2}+\eta\|f\|^{2}_{H_{K}}\\ =\|\left(Y-\beta^{\top}\boldsymbol{K}\right)A\|^{2}_{F}+\eta tr\left(\beta^{\top}\boldsymbol{K}\beta\right),\\ \end{split}$ (19)
where $\boldsymbol{K}\in\mathbb{R}^{(n+m)\times(n+m)}$ is the kernel matrix, with $\boldsymbol{K}_{ij}=\boldsymbol{K}(z_{i},z_{j})$. The label information of the source and target data is denoted by $Y=[y_{1},...,y_{n+m}]$. For the target data, the corresponding label information is the pseudo label generated by the learned classifier $f(\cdot)$. $tr(\cdot)$ is the trace operation.
For the second part in (16), through applying (17), we can represent the calculation of dynamic distribution adaptation (12) as
$\overline{D}_{f}\left(T_{S},T_{T}\right)=tr\left(\beta^{\top}\boldsymbol{K}M\boldsymbol{K}\beta\right),$ (20)
where $M=(1-\mu)M_{0}+\mu\sum^{C}_{c=1}M_{c}$. $c$ is the class category, and $C$ refers to the total number of classes. $M_{0}$ and $M_{c}$ are defined as
$\left(M_{0}\right)_{ij}=\begin{cases}\frac{1}{n^{2}}&z_{i},z_{j}\in T_{S}\\ \frac{1}{m^{2}}&z_{i},z_{j}\in T_{T}\\ -\frac{1}{mn}&otherwise\end{cases},$ (21)
$\left(M_{c}\right)_{ij}=\begin{cases}\frac{1}{n^{2}_{c}}&z_{i},z_{j}\in T_{S}^{(c)}\\ \frac{1}{m^{2}_{c}}&z_{i},z_{j}\in T_{T}^{(c)}\\ -\frac{1}{m_{c}n_{c}}&\begin{cases}z_{i}\in T_{S}^{(c)},z_{j}\in T_{T}^{(c)}\\ z_{i}\in T_{T}^{(c)},z_{j}\in T_{S}^{(c)}\end{cases}\\ 0&otherwise\end{cases}.$ (22)
Here, $n_{c}=|T_{S}^{(c)}|$ and $m_{c}=|T_{T}^{(c)}|$. The third part in (16) (the Laplacian regularization term $R_{f}\left(T_{S},T_{T}\right)$) could be computed as
$\begin{split}R_{f}\left(T_{S},T_{T}\right)=\sum^{n+m}_{i,j=1}W_{ij}\left(f\left(z_{i}\right)-f\left(z_{j}\right)\right)^{2}\\ =\sum^{n+m}_{i,j=1}f\left(z_{i}\right)L_{ij}f\left(z_{j}\right),\\ \end{split}$ (23)
where $W_{ij}$ is a pairwise similarity matrix, and $L=D-W$ is the normalized graph Laplacian matrix. $D$ is a diagonal matrix, with $D_{ii}=\sum^{n+m}_{j=1}W_{ij}$. $W$ could be represented as
$W_{ij}=\begin{cases}sim\left(z_{i},z_{j}\right)&z_{i}\in N_{p}\left(z_{j}\right)orz_{j}\in N_{p}\left(z_{i}\right),\\ 0&otherwise,\end{cases}$ (24)
where $sim(\cdot,\cdot)$ is a similarity function (e.g., cosine distance) used to assess the proximity between two points. $N_{p}(z_{i})$ represents a set of $p$ nearest points to $z_{i}$. $p$ is a hyperparameter.
Similarly, by applying (17), the Laplacian regularization given in (23) can be expressed as
$R_{f}\left(T_{S},T_{T}\right)=tr\left(\beta^{\top}\boldsymbol{K}L\boldsymbol{K}\beta\right).$ (25)
Then, by substituting (19), (20) and (25)into the (16), the classifier $f$ can be calculated as:
$\begin{split}f=\mathop{\arg\min}_{f\in H_{K}}\|\left(Y-\beta^{\top}\boldsymbol{K}\right)A\|^{2}_{F}+\eta tr\left(\beta^{\top}\boldsymbol{K}\beta\right)\\ +tr\left(\beta^{\top}\boldsymbol{K}\left(\lambda M+\rho L\right)\boldsymbol{K}\beta\right).\end{split}$ (26)
Setting $\frac{\partial f}{\partial\beta}=0$, (26) could be solved as:
$\begin{split}\beta^{*}=\left(\left(A+\lambda M+\rho L\right)\boldsymbol{K}+\eta I\right)^{-1}AY^{\top}.\end{split}$ (27)
After obtaining $\beta^{*}$, the classifier $f$ can be determined using (26). Then, the corresponding soft label of the target domain at $\iota$th iteration could be denoted as $\hat{y}_{t}^{\iota}=f^{\iota}(z_{t})$. In the implementation, through $l$ iterations, the adaptive factor $\mu$ in the dynamic distribution alignment is continuously updated, and the classifier learning leverages the corresponding updated $\mu$ to further refine the pseudo label of the target domain.
Algorithm 1 The algorithm flow of the proposed M3D model.
1:- Data matrix $X=\left[X_{s},X_{t}\right]$, source labels $y_{s}$, subspace dimension $d$, regularization parameters $\lambda,\beta,\rho$ and the iteration number $\iota$;
2:Classifier $f(\cdot)$;
3:# Manifold Feature Transformation
4:Learn the manifold feature transformation kernel $G$ and obtain the manifold feature representation $z$ using (8);
5:Train a weak classifier based on the source data $D_{S}$, and get initialized soft labels $\hat{y}_{t}^{0}$ for the target data $D_{T}$;
6:Construct the geodesic line kernel $K$ using the transformed features $z_{s}=z_{1:n}$ and $z_{t}=z_{n+1:n+m}$;
7:for $\iota=1$ to $l$ do
8:# Dynamic Distribution Alignment
9:     Calculate the adaptive factor $\mu$ using (15), and compute $M_{0}$ and $M_{c}$ using (21) and (22);
10:# Classifier Learning
11:     Solve the equation using (27) to compute $\beta^{*}$ and obtain the classifier $f$ using (26);
12:     Update the soft labels for the target data $\hat{y}_{t}^{\iota}=f(z_{t})$;
13:end for
14:Return the classifier $f$;
15:# Ensemble Learning
16:Perform ensemble learning based on the obtained soft labels in the loop $\{\hat{y}_{t}^{\iota}\}_{\iota=1}^{l}$.
III-D Ensemble Learning
Ensemble learning with LinkCluE algorithm [36] is adopted here to strategically reduce classification bias inherent in individual models and improve the overall generalization performance of the model. Based on the obtained $\hat{y}_{t}$ in a certain round of classifier learning, the ensemble learning leverages the strengths over a group of $\hat{y}_{t}^{\iota}$ ($\iota=1,\dots,l$. Empirically, $l$ is set to 10) and form together to generate a strong prediction results $\hat{y}_{t}$.
Fig. 2 illustrates the two main steps involved in the LinkCluE algorithm: the link-based similarity matrix function and the consensus function. For the link-based similarity matrix function, it constructs a similarity matrix based on the links or relationships between data points in terms of Connected-Triple-Based Similarity (CTS), SimRank-Based Similarity (SRS), and Approximate SimRank-Based Similarity (ASRS). This matrix serves as a foundation for understanding the complex, often non-linear relationships that exist between data points, leveraging the structure of the data to enhance the learning process. For the consensus function, it integrates the outcomes of various individual models within the ensemble by leveraging three different hierarchical agglomerative clustering algorithms, Single Linkage (SL), Complete Linkage (CL), and Average Linkage (AL). This step is crucial for mitigating individual model biases and errors, leading to more accurate and generalizable results.
The overall algorithm of the proposed M3D is illustrated in Algorithm 1.
Figure 2: The main process of the LinkCluE ensemble algorithm. (1) Link-based Similarity Matrix Function: takes a set of input vectors $\hat{y}_{t}^{\iota}$ ($\iota=1,\dots,l$) and constructs a similarity matrix as its output. (2) Hierachical Clustering: processes the similarity matrix through a hierarchical clustering algorithm to produce the final clustering result, denoted as $\hat{y}_{t}$.
IV Experimental Results
IV-A Emotional Databases
The model performance is carefully validated on two well-known publicly available databases: SEED [11] , SEED-IV [37]. The SEED database includes EEG emotion data collected from 15 participants. Each participant was exposed to a total of 15 movie clips designed to evoke three types of emotions (positive, negative, and neutral). The SEED-IV database also includes EEG emotion data from 15 participants, exposed to a total of 24 movie clips aimed at eliciting four types of emotions (happy, sad, neutral, and fear). For both SEED and SEED-IV databases, participants underwent three sessions spread across separate days, with a one-week break between sessions. The EEG signals were recorded using the ESI Neuroscan system with 62 channels.
The EEG preprocessing followed established procedures in the literature. The process began with the downsampling of EEG signals to 200 Hz, followed by the removal of artifacts such as the Electrooculogram (EOG) and Electromyography (EMG). A bandpass filter with a 0.3-50 Hz range was then applied to improve the signal quality. Following this, each trial was segmented into multiple 1-second data samples. To extract emotional related information, DE features were extracted across five frequency bands: Delta (1-3 Hz), Theta (4-7 Hz), Alpha (8-13 Hz), Beta (14-30 Hz), and Gamma (31-50 Hz). Thus, for each 1-second data sample, a total of 310 features (5 frequency bands $\times$ 62 channels) were characterized, which serves as the input for the model.
To further assess the generalizability of the proposed M3D model beyond controlled experimental conditions, we extend model validation to a real-world clinical EEG database obtained from Hospital Universiti Sains Malaysia (HUSM). This dataset consists of EEG recordings from 27 healthy individuals (mean age: 38.28±15.64 years) and 29 individuals diagnosed with MDD (Major Depressive Disorder) (mean age: 40.33±12.86 years). Participant classification adhered to the international diagnostic criteria outlined in the Diagnostic and Statistical Manual-IV (DSM-IV). This dataset used different devices and settings compared to the SEED and SEED-IV databases, with EEG recordings collected using the Brain Master Discovery amplifier, 19 electrodes, and a sampling rate of 256 Hz. Preprocessing involved bandpass filtering (0.5-70 Hz) and notch filtering at 50 Hz to eliminate power-line noise. The experimental protocol included EEG recordings under both open-eye (EO) and closed-eye (EC) conditions, each lasting 5 minutes. During the EO condition, participants were instructed to remain relaxed while minimizing eye movements to maintain data quality.
IV-B Experimental Protocols
In order to thoroughly assess the robustness and stability of the proposed model and ensure a detailed comparison with existing literature, two types of experimental protocols are conducted.
  • Cross-subject single-session leave-one-subject-out cross-validation. The data samples from N subjects (SEED/SEED-IV: N=14) in the first session are used as the source domain, and the data samples from the remaining 1 subject in the same session are used as the target domain. This procedure repeats N+1 times (SEED/SEED-IV: N=14), ensuring that each participant is treated as the target domain at least once.
  • Cross-subject cross-session leave-one-subject-out cross-validation. The data samples from N subjects (SEED/SEED-IV: N=14) across all three sessions are used as the source domain, and the data samples from the remaining 1 subject across all three sessions are used as the target domain. This procedure also repeats N+1 times (SEED/SEED-IV: N=14), ensuring that each participant is treated as the target domain at least once.
In the present study, we adopt six widely recognized machine-learning classifiers to generate the initial classification results ($\hat{y}_{t}^{0}$) as mentioned in Section III-C. The adopted machine-learning classifiers include KNN[38], SVM[39], DT[40], Adaboost classifier[41], GNB classifier[42], and Bagging classifier[43]. In the SVM classifier, the radial basis function (RBF) kernel is utilized. In ensemble learning, the application of three similarity matrix functions (CTS, SRS, and ASRS) combined with three hierarchical agglomerative clustering algorithms (SL, CL, and AL) yields a total of 9 possible combinations. Here, the combination that achieves the highest accuracy in predicting target domain labels will be chosen as the optimal outcome. Notably, all the best results are achieved using the DT classifier and the CTS-SL approach. For performance evaluation, a comprehensive analysis of the model’s efficacy is conducted using the accuracy and F1-score metrics.
IV-C Experimental Results on Cross-Subject Single-Session
The experimental results on SEED and SEED-IV database under the cross-subject single-session cross-validation are reported in Table I. For the SEED database, the best performance is obtained when DT is adopted as the initial classifier, where the corresponding accuracy and F1-score are 84.57% and 76.11%. We compare the obtained experimental results with the existing literature, as shown in Table II. The comparison results show that the proposed non-deep transfer learning model (M3D) demonstrates superior performance. In comparison to the best results obtained by the traditional machine learning methods, our model shows a performance enhancement of 8.26%. Furthermore, it performs comparably or even surpasses some deep learning methods, such as PGCN [44], ADA [15], TANN [45] and DG-DANN [46].
TABLE I: Cross-subject single-session leave-one-subject-out cross-validation on SEED and SEED-IV with different initial classifiers.
Database KNN SVM DT Adaboost GNB Bagging
SEED Accuracy 84.46 84.51 84.57 84.10 84.10 84.10
F1-score 75.77 75.84 76.11 74.49 74.49 74.49
SEED-IV Accuracy 58.25 58.82 60.94 58.89 58.42 58.25
F1-score 48.80 47.51 50.49 48.06 45.97 48.80
TABLE II: Performance comparison on SEED using cross-subject single-session leave-one-subject-out cross-validation. The second-best traditional machine learning methods result is underlined.
Methods $P_{acc}$ Methods $P_{acc}$
Traditional machine learning methods
SVM[13] 58.18/13.85 TKL[47] 63.54/15.47
TCA[13] 64.00/14.66 SCA[46] 66.33/10.60
SA[48] 69.00/10.89 KPCA[13] 69.02/09.25
DICA[46] 69.40/07.80 GFK[47] 71.31/14.09
T-SVM[47] 72.53/14.00 TPT[47] 76.31/15.89
Deep learning methods
MLP[13] 61.01/12.38 DNN[13] 61.01/12.38
DCORAL[49] 62.14/07.98 GELM[50] 70.80/00.74
DAN[49] 73.54/06.09 DDC[49] 74.34/09.05
R2G-STLT[51] 77.96/06.38 DANN[13] 79.79/13.14
MS-MDA[49] 79.67/08.01 DGCNN[52] 79.95/09.02
MMD [15] 80.88/10.10 SimNet[53] 81.58/05.11
GECNN[54] 82.46/10.83 Bi-DANN[47] 83.28/09.60
DGGN[50] 83.84/10.26 Saliency[55] 84.11/02.90
BiDANN-S[47] 84.14/06.87 R2G-STNN[56] 84.16/07.63
DG-DANN[46] 84.30/08.32 TANN[45] 84.41/08.75
ADA[15] 84.47/10.65 PGCN[44] 84.59/08.68
BiHDM[48] 85.40/07.53 ST-SCGNN[57] 85.90/04.90
M3D 84.57/09.49
For the SEED-IV database, the best accuracy is obtained when DT is used as the initial classifier, where the corresponding accuracy and F1-score are 60.94% and 50.49%, respectively. The performance comparison results with the existing methods are reported in Table III. Compared to the best results reported in the traditional machine learning methods, the proposed M3D demonstrates a slight performance advantage of 0.72%. Furthermore, it performs comparably or even surpasses some deep learning methods, including ScalingNet [51], ROADAN [58], WGAN-GP [59] and MS-ADA [49].
TABLE III: Performance comparison on SEED-IV using cross-subject single-session leave-one-subject-out cross-validation. The second-best traditional machine learning methods result is underlined.
Methods $P_{acc}$ Methods $P_{acc}$
Traditional machine learning methods
TCA[60] 37.01/10.47 GFK[60] 44.04/09.31
GFHF[60] 49.29/07.60 JDA[60] 54.57/05.47
GAKT[60] 58.49/06.23 MIDA[60] 60.22/08.69
Deep learning methods
DCORAL[49] 40.50/10.05 DAN [49] 42.03/05.66
DDC[49] 44.06/06.01 JD-IRT[61] 52.12/15.88
RGNN[62] 53.49/10.57 JAGP[62] 54.37/19.49
DANN[63] 54.63/08.03 GECNN[55] 55.01/10.29
MS-MDA[64] 59.34/05.48 MS-ADA[49] 59.29/13.65
WGAN-GP[59] 60.60/15.76 ROADAN[58] 60.75/10.40
ScalingNet[51] 60.75/10.40 MFA-LR[62] 64.96/14.04
M3D 60.94/08.84
IV-D Experimental Results on Cross-Subject Cross-Session
To evaluate the efficiency and stability of the proposed M3D in managing variations in EEG data collected from the same subject at different times, we also conduct a model validation under the cross-subject cross-session leave-one-subject-out cross-validation protocol. The experimental results are reported in Table IV, when different initial classifiers are used. The best model performance of SEED database is achieved when DT is adopted as the initial classifier, where the accuracy and F1-score are 77.05% and 66.46%, respectively. The performance results against the existing methods are compared in Table V. It shows, the proposed model demonstrates an accuracy improvement of 5.18% in comparison to traditional machine learning methods. Further, it exhibits performance that is comparable, to some extent, with deep learning methods, showing the effectiveness of the proposed M3D in managing variations in data collected from the same subject across different time periods.
TABLE IV: Cross-subject cross-session leave-one-subject-out cross-validation on SEED and SEED-IV with different initial classifiers
Database KNN SVM DT Adaboost GNB Bagging
SEED Accuracy 74.79 75.98 77.05 76.60 76.44 76.47
F1-score 60.88 64.67 66.46 65.43 65.28 65.30
SEED-IV Accuracy 64.55 64.55 64.90 64.71 64.71 64.71
F1-score 53.09 53.09 52.20 53.20 53.20 53.20
TABLE V: Performance comparison on SEED using cross-subject cross-session leave-one-subject-out cross-validation. The second-best traditional machine learning methods result is underlined.
Methods $P_{acc}$ Methods $P_{acc}$
Traditional machine learning methods
KNN[65] 60.66/07.93 SA[26] 61.41/09.75
TCA[17] 64.02/07.96 GFK[28] 66.02/07.59
SVM[66] 68.15/07.38 CORAL[67] 68.15/07.83
RF[68] 69.60/07.64 Adaboost[69] 71.87/05.70
Deep learning methods
TSception[70] 64.30/09.80 CDCN[70] 69.30/09.40
STEENN[70] 72.00/09.00 A-LSTM[71] 72.18/10.85
STRNN[70] 73.20/11.10 3D-CNN[70] 74.20/07.80
MMResLSTM[70] 74.40/10.40 ACRNN[70] 76.30/08.10
MetaEmotionNet[70] 77.50/08.80 DGCNN[72] 79.95/09.02
DAN [73] 81.04/05.32 DANN[74] 81.08/05.88
DCORAL[75] 81.97/05.16 DDC[76] 82.17/04.96
M3D 77.05/07.61
For the SEED-IV database, the performance across various initial classifiers is relatively consistent, with the highest accuracy reaching 64.90% when DT is employed as the initial classifier. Table VI reports the performance comparison results against the existing machine learning and deep learning methods. Compared to the best results reported in various literature on machine learning methods, the proposed M3D shows a performance improvement of 0.46%. Furthermore, it performs comparably or even outperforms some deep learning methods, including MetaEmotionNet [70], MS-ADA [49] and DAN [13].
TABLE VI: Performance comparison on SEED-IV using cross-subject cross-session leave-one-subject-out cross-validation. The second-best traditional machine learning methods result is underlined.
Methods $P_{acc}$ Methods $P_{acc}$
Traditional machine learning methods
KNN[65] 40.83/07.28 GFK[28] 45.89/08.27
CORAL[67] 49.44/09.09 RF[68] 50.98/09.20
SVM[13] 51.78/12.85 KPCA[13] 51.76/12.89
TPT[13] 52.43/14.43 Adaboost[69] 53.44/09.12
TCA[13] 53.97/08.05 SA[48] 64.44/09.47
Deep learning methods
DCORAL[49] 40.50/10.05 DAN[49] 42.03/05.66
DDC[49] 44.06/06.01 ACRNN[70] 49.20/09.20
DNN[13] 49.35/09.74 MLP[13] 49.35/09.74
MMResLSTM[70] 51.10/11.40 DGCNN[52] 52.82/09.23
STRNN[70] 53.20/07.40 3D-CNN[70] 54.10/10.90
CDCN[70] 54.50/13.10 DANN[13] 54.63/08.03
A-LSTM[71] 55.03/09.28 TSception[70] 56.20/09.50
STFFNN[70] 56.70/07.30 MS-MDA[49] 57.92/10.12
DAN [13] 58.87/08.13 MS-ADA[49] 59.29/13.65
MetaEmotionNet[70] 61.20/08.30 Bi-DANN[47] 65.59/10.39
M3D 64.90/10.23
IV-E Experimental Results on clinical EEG database
To assess the generalizability of the proposed M3D model in real-world clinical applications, we evaluate its performance on the MDD database (n=56). Following previous work [77], we adopt the cross-subject single-session 10-fold cross-validation protocol for fair comparison. The validation results are presented in Table VII, demonstrating consistent performance across various initial classifiers. In particular, the highest accuracy is achieved using DT as the initial classifier, reaching 82.16% in the EC session and 78.49% in the EO session. For comparative analysis, we report performance differences against existing methods in Table VIII. The proposed M3D model outperforms the best-performing traditional machine learning approach by 7.95% and 4.28% in the EC and EO sessions, respectively. Among deep learning methods, DCORAL [75] achieves the highest reported accuracy in prior studies, with 70.88±10.53% in the EC session and 74.49±08.17% in the EO session. M3D surpasses these results by 11.28% and 4% in the EC and EO sessions, respectively, demonstrating its superior ability to capture the complex patterns associated with MDD. These results underscore the effectiveness of M3D in real-world clinical settings, establishing it as a promising tool for practical clinical applications.
TABLE VII: Cross-subject single-session 10 fold cross-validation on MDD with different initial classifiers.
Session KNN SVM DT Adaboost GNB Bagging
EC Accuracy 82.16 82.16 82.16 82.16 80.37 80.37
F1-score 81.89 81.89 81.89 81.89 80.43 80.43
EO Accuracy 74.92 74.92 78.49 74.92 74.92 74.92
F1-score 73.97 73.97 77.67 73.97 73.97 73.97
TABLE VIII: Performance comparison on MDD using cross-subject 10-fold cross-validation. Models reproduced by the authors are indicated with an asterisk (*), while the second-best traditional machine learning method’s results are underlined. EC: closed-eye; EO: open-eye.
      Methods       EC session       EO session
      $P_{acc}$       $P_{acc}$
      Traditional machine learning methods
      SVM*[39]       35.56/10.59       35.56/10.59
      CORAL*[67]       60.27/15.25       60.12/13.58
      RF*[68]       69.44/19.57       69.44/19.57
      Adaboost*[41]       70.49/16.70       71.00/19.62
      KNN*[65]       72.25/18.32       72.25/18.32
      SA*[26]       74.21/11.98       74.21/11.98
      Deep learning methods
      DAN* [13]       64.19/05.30       64.99/09.01
      DDC*[76]       64.65/04.84       65.69/08.08
      DANN*[13]       66.49/11.08       70.25/09.67
      DCORAL*[75]       70.88/10.53       74.49/08.17
      M3D       82.16/11.35       78.49/13.79
TABLE IX: Performance comparison without classifier learning and ensemble learning, with only classifier learning, and with both classifier learning and ensemble learning.
Manifold Feature Transformation Dynamic Distribution Alignment and Classifier Learning Ensemble Learning Accuracy F1-Score 48.27/14.39 35.42 69.51/16.40 56.53 71.01/13.42 59.41 79.66/09.34 69.94 84.57/09.49 76.11
V Discussion
V-A Ablation Study
To comprehensively examine the contributions of the model components within the proposed M3D model, we conduct a thoroughly ablation study. Specifically, we design four model variations to isolate the effects of key modules. (1) Manifold Feature Transformation Only: the initial decision tree classifier is trained to evaluate model performance after applying only the manifold feature transformation module. (2) Dynamic Distribution Alignment and Classifier Learning Only: the DE feature is used as input, focusing solely on dynamic distribution alignment and classifier learning. (3) Without Manifold Feature Transformation: the DE feature serves as input for dynamic distribution alignment, classifier learning, and ensemble learning, excluding the manifold feature transformation module. (4) Without Ensemble Learning: the ensemble learning module is removed to evaluate its impact on model performance. Note here that all the presented experimental results in this section are based on the SEED database using cross-subject single-session leave-one-out cross-validation. The decision tree is employed as the initial classifier, with 10 iterative loops for evaluation.
Table IX presents the results of the ablation study, demonstrating that the proposed M3D model achieves the best performance by effectively integrating all modules. (1) Manifold Feature Transformation Only: When only the manifold feature transformation module is used, the model performance is the lowest, with a classification accuracy of 44.73±14.39%. (2) Dynamic Distribution Alignment and Classifier Learning Only: Using only dynamic distribution alignment and classifier learning yields better results than using only manifold feature transformation, achieving an accuracy of 69.51±16.40%. (3) Without Manifold Feature Transformation: Adding ensemble learning to dynamic distribution alignment and classifier learning further improves accuracy to 71.01±13.42%, reflecting a 1.5% improvement. This highlights the effectiveness of ensemble learning in enhancing EEG-based emotion recognition performance. However, compared to the full model, removing the manifold feature transformation module results in a significant accuracy drop from 84.57±09.49% to 71.01±13.42%. This underscores the importance of feature transformation before distribution alignment in effectively reducing domain differences and improving model performance. (4) Without Ensemble Learning: Removing the ensemble learning module negatively impacts the model, leading to a reduced accuracy of 79.66±09.34%. Additionally, features in the Grassmann manifold space prove more robust than those in the original feature space, improving distribution alignment. All four ablation models differ significantly from the full model (p $<$ 0.0001, paired t-test, FDR corrected).
TABLE X: Performance comparison using TCA and PCA.
         Methods          PCA          TCA
         Accuracy          52.93          84.57
         F1-Score          50.55          76.11
TABLE XI: Performance comparison with different $\mu$ values.
$\mu$ 0 0.5 1 dynamic
Accuracy 78.57 81.55 80.31 84.57
F1-Score 69.68 72.56 71.28 76.11
V-B Investigating factors affecting model performance in different modules
To evaluate the efficacy of TCA in the manifold feature transformation, we assess the model performance by comparing the use of TCA against substituting TCA with Principal Component Analysis (PCA). As shown in Table X, the model exhibits significantly better performance in terms of accuracy and F1-Score when comparing TCA with PCA. This improvement can be attributed to TCA’s ability to enhance model performance by simultaneously mapping and reducing dimensions, and by maximizing the similarity between datasets from two domains. It would facilitate more effective dynamic distribution alignment and classifier categorization.
TABLE XII: Performance comparison with different ensemble learning.
Methods Last Averaging Voting LinkCluE
Accuracy 79.66 78.60 80.62 84.57
F1-Score 69.94 66.97 71.59 76.11
TABLE XIII: The emotion recognition accuracy (%) without and with ensemble learning in iterative learning process.
Iteration 2 5 10 20 30 50
w/o ensemble learning 72.85 75.09 74.61 74.09 74.37 73.45
with ensemble leaning 80.39 83.81 84.57 83.84 83.34 82.55
To evaluate the effectiveness of dynamic distribution alignment, we opted for a fixed $\mu$ value approach, rather than using an adaptive $\mu$. The $\mu$ values are set at 00, $0.5$, and $1$. At $\mu=0$, the model exclusively focuses on the marginal distribution. With $\mu=0.5$, both marginal and conditional distributions are considered equally. At $\mu=1$, the model solely emphasizes the conditional distribution. As reported in Table XI, it shows that employing an adaptive $\mu$ can significantly enhance model performance, resulting in a 4.26% improvement in accuracy. The results demonstrate that dynamic distribution alignment can effectively quantify the disparity between source and target domains and accurately evaluate the significance of the differences between marginal and conditional distributions. This capability is particularly advantageous for the classification performance, which is tailored for addressing variations between distributions.
To evaluate the impact of ensemble learning on model performance, we adjust our approach by incorporating various ensemble learning techniques, namely the averaging method, the voting method, and the LinkCLuE method. Note here that LinkCLuE is the method implemented in the proposed M3D. Based on the obtained classification results in the classifier learning ($\hat{y}_{t}^{(\iota)}$ ($\iota=1,\cdots,10$)), the effectiveness of ensemble learning is analyzed. Furthermore, we drew comparisons to scenarios devoid of ensemble learning strategies by directly employing the final classification outcome from the last iteration loop in the classifier learning, denoted as $\hat{y}_{t}^{(\iota)}$ (with $\iota=10$), which serves as our baseline. The experimental comparison results are reported in Table XII. It shows that ensemble learning with voting and LinkCluE could enhance the reliability of classification results and improve the overall performance. This confirms the effectiveness of ensemble algorithms. Furthermore, the LinkCLuE method outperforms the voting method, demonstrating a superior ability to synthesize classification outcomes across iterations and to discern the underlying patterns of the classification results.
V-C Parameter effect
To examine the classifier learning process, we also investigate how sensitive the model is to changes in the number of iterations ($\iota$) in the learning process. Here, we vary $\iota$ value in a range of $[2,5,10,20,30,50]$, and present the corresponding learning results in Table XIII. It shows the model’s classification performance experiences fluctuations during the initial few iterations but achieves stability beyond $\iota=10$ iterations. This demonstrates the training advantage of M3D in cross-domain tasks.
VI Conclusion
In this paper, we propose a novel non-deep transfer learning framework, termed M3D, for cross-subject and cross-session EEG-based emotion recognition, as well as clinical mental disorder recognition. The framework employs a manifold-based domain adaptation with dynamic distribution alignment, adaptively balancing marginal and conditional distributions in the Grassmann manifold space. This enhances intrinsic data representation and reduces domain disparity. Additionally, optimized classifier learning with ensemble techniques improves robustness and reliability in handling EEG data variations. Extensive experiments on two benchmark databases validate the framework under two evaluation protocols. Furthermore, we demonstrate the model’s effectiveness on a real clinical MDD dataset. Compared to existing methods, M3D achieves promising results against non-deep approaches and performance comparable to deep learning models, with an average improvement of 4.47% in classification accuracy. These findings highlight the potential of non-deep transfer learning in addressing individual and session variations in affective brain-computer interfaces. On the other hand, in the dynamic distribution alignment module, we did not impose constraints on the adaptive factor $\mu$ during the iteration process, which may lead to suboptimal feature alignment in certain cases. Addressing this limitation in future work could further enhance the stability and generalizability of the framework.